OpenStudy (anonymous):

slope -4;y-intercept (0,-3/5)

OpenStudy (saifoo.khan):

y = -4x -3/5

OpenStudy (anonymous):

To do this, do y-y1 = m(x-x1)

OpenStudy (anonymous):

What this means is: y = just regular old y y1 = the y value from your point (-3/5) x = just regular x x1 = 0

OpenStudy (anonymous):

so, y-(3/5) = -4(x-0)

OpenStudy (anonymous):

Sorry... y-(NEGATIVE 3/5) = -4(x-0)

OpenStudy (anonymous):

you get: y+ 3/5 = -4x + 0 Move the 3/5 to the right and you have: y = -4x - 3/5 Make sense? ^_^

OpenStudy (anonymous):

yes and thanks guys